Javier invests money in two simple interest accounts. He invests twice as much in an account paying 13% as he does in an account paying 6%. If he earns $104.00 in interest in one year from both accounts combined, how much did he invest altogether?
Total Principal in Both Accounts = $

Found 2 solutions by JThomson, addingup:
Answer by JThomson(12)


Let the amount invested (Principal) for bank B be y;

Interest on Bank B with interest rate at 6% is;

(y×6×1)/100

Similarly amount invested for bank A is 2y;

Interest on bank A is given by;

(2y×13×1)/100

Total Interest is 104.00

(y×6×1)/100 + (2y×13×1)/100 = 104

6y+26y = 10400

32y = 10400

y = $325.00 ; This is the interest on "Bank B"

2y = 2×325 = $650.00; This is the interest on "Bank A".

Total amount invested is 325+650 = $975

Answer by addingup(3677)

0.06x + 0.13(2x) = 104
0.06x + 0.26x = 104
0.32x = 104
x = 325 this is the amount at 6%
2x = 2(325) = 650 this is the amount at 13%
------------------------------------
Check:
0.06*325 = 19.5
0.13*650 = 84.5
. . . . .-------
Total . . 104
The answer is correct